OpenStudy (anonymous):

f(x)=3x^2 +5 and g(x)= 7x-2 Verify g(x+2) notequalto g(x)+g(2)

OpenStudy (jdoe0001):

\(\large { f(x)=3x^2 +5 \qquad g(x)= 7x-2 \\ \quad \\ \begin{cases} g({\color{brown}{ x+2}})=7({\color{brown}{ x+2}})-2\implies &? \\ \quad \\ g({\color{brown}{ 2}})=7({\color{brown}{ 2}})-2\implies &? \end{cases} }\)